OpenStudy (anonymous):

What is the slope and y-intercept of the line represented by this equation? 5x– 2y = 6

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So we need to put this equation in y = mx + b form. 5x - 2y = 6 --- subtract 5x from both sides -2y = -5x + 6 --- now divide both sides by -2 y = 5/2x - 3 so your slope (m) = 5/2 and your y intercept (b) = -3